OpenStudy (anonymous):

Prove the identity (cos 3x+cos x)/(sin3x-sinx)=cotx

OpenStudy (mertsj):

\[\cos x+\cos y=2\cos \frac{x+y}{2}\cos \frac{x-y}{2}\]

OpenStudy (mertsj):

So cos 3x+cosx=2cos2xcosx Also: \[\sin x+\sin y=2\cos \frac{x+y}{2}\cos \frac{x-y}{2}\]

OpenStudy (mertsj):

So sin3x-sinx = 2cos2xsinx

OpenStudy (mertsj):

So \[\frac{\cos 3x+\cos x}{\sin 3x-\sin x}=\frac{2\cos 2x \cos x}{2\cos 2x \sin x}=\frac{\cos x}{\sin x}=\cot x\]

OpenStudy (anonymous):

\[\sin x+\sin y=2\sin \frac{x+y}{2}\cos \frac{x-y}{2} ****\]